OpenStudy (anonymous):

there is a mixture of 20% red paint, 35% yellow paint, and 45% water. If 5 quarts of red paint are added to 20 quarts of the mix, what is the percentage of red paint in the new mixture?

OpenStudy (tkhunny):

How many total quarts in the original mixture?

OpenStudy (anonymous):

20 quarts?

OpenStudy (triciaal):

20% of 20 = 4 quarts of red originally add 5 quarts now you have 9 quarts of red in the total mix of (20 + 5 ) as a percentage = 9/25* 100 = 36 %

OpenStudy (triciaal):

check 9/25 + 7/25 + 9/25 = 25/25 red , yellow and water respectively 36% + 28% + 36% = 100%